OpenStudy (anonymous):

Need help! Solve. 4y + 16 > –8 A. y > –6 B. y > –2 C. y > 2 D. y > 6 I got y < -6

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it looks like you switched the inequality for some reason

OpenStudy (anonymous):

4y + 16 > -8 Subtract 16 from each side: 4y > -24 Divide each side by 4: y > -6

OpenStudy (anonymous):

Yeah because it has a negative number

OpenStudy (anonymous):

change the inequality ONLY IF you are multiplying or dividing by a negative number

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

it does not matter whether the solution is positive or negative it only matters if you multiply or divide by a negative number

OpenStudy (anonymous):

wait I am.

OpenStudy (anonymous):

since at the last step you are dividing by 4, positive 4, you don't switch anything

OpenStudy (anonymous):

4 divided by -24 changes something

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

4y + 16 > –8 subtract 16 from both sides -8-16=?

OpenStudy (anonymous):

-8-16= -24

pooja195 (pooja195):

Good now we have 4y>-24 divide both sides by 4 to isolate y 4/4=? -24/4=?

OpenStudy (anonymous):

-6

pooja195 (pooja195):

so the answer would be...?

OpenStudy (anonymous):

y<-6

pooja195 (pooja195):

why did you flip the sign?

OpenStudy (anonymous):

because I divided it with a negative sign

pooja195 (pooja195):

We only flip the sign when we divide BY a negative so if it was -4 we would flip the sign

OpenStudy (anonymous):

so it matters which order we're dividing it?

pooja195 (pooja195):

no it matters what you are dividing by not the order if i asked you to divide both sides by -4 y<-6 would be corrct but we are dividing by 4 only

OpenStudy (anonymous):

ok